OpenStudy (anonymous):

what is 1.25 as a fraction plzz help ill give medal and fan

OpenStudy (anonymous):

@mathmath333 plzzz help

OpenStudy (anonymous):

@magepker728 help plzzz

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{1.25\\~\\ =\dfrac{125}{100}\\~\\ =\dfrac{25\times 5}{25\times 4}\\~\\ =\dfrac{ 5}{4}\\~\\ }\end{align}\)

OpenStudy (anonymous):

@magepker728 could u explain that better

OpenStudy (anonymous):

isnt the answer one an one fourth

OpenStudy (mathmath333):

yes it is but that is a mixed fraction

OpenStudy (anonymous):

o

OpenStudy (anonymous):

\[1 = 4/4, .25 = 1/4.... 4/4 + 1/4 = 5/4\]

OpenStudy (anonymous):

ok

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ =\dfrac{ 5}{4}\\~\\ =\dfrac{ 4+1}{4}\\~\\ =1+\dfrac{ 1}{4}\\~\\ =1\cfrac{ 1}{4} }\end{align}\)

OpenStudy (anonymous):

kk got it thx